Circular No. 7430

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 2000 K1

F. Shelly, Lincoln Laboratory, Massachusetts Institute of Technology, reports the discovery by LINEAR of a comet. Additional CCD observations have also been received following posting on the NEO Confirmation Page:

2000 UT α_{2000} δ_{2000} m_{1} Observer
May 26.31260 16^{h}35^{m}59.^{s}91 +9^{o}43′34′′.2 18.0 LINEAR
26.37246 16 35 56.66 +9 44 00.8 17.7  _{”}
26.84775 16 35 32.53 +9 47 31.6 16.0 Tichá
26.85382 16 35 32.14 +9 47 34.5  _{”}
26.90561 16 35 29.49 +9 47 56.8  Kušnirák
26.91633 16 35 28.94 +9 48 01.4  _{”}
26.93501 16 35 27.98 +9 48 10.0  Galád
26.93671 16 35 27.89 +9 48 10.7 14.9  _{”}

M. Blythe, F. Shelly, M. Bezpalko, M. Elowitz, and R. Manguso (Lincoln Laboratory ETS). 1.0-m f/2.15 reflector. Measurers J. Stuart, R. Sayer, and J. B. Evans. Tail in p.a. 160^{o}.

J. Tichá and M. Tichý (Kleť). 0.57-m f/5.2 reflector. Coma diameter ∼

12′′ and wide 20′′ tail in p.a. 205^{o}. P. Kušnirák (Ondřejov). 0.65-m f/3.6 refl. Broad tail ∼ 2′ long in p.a. 140^{o}. A. Galád and J. Tóth (Modra). 0.6-m f/5.5 reflector. Tail in p.a. ∼ 160^{o}.

COMET 2000 K2

Shelly reports the discovery by LINEAR of another comet showing a tail in p.a. 240^{o}. The additional CCD observations below were reported following posting on the NEO Confirmation Page (observer details given above). Tichá reports the comet to be slightly diffuse, and Kušnirák reports a faint tail ∼ 20′′ long in p.a. 220^{o}; Galád reported cometary appearance.

2000 UT α_{2000} δ_{2000} m_{1} Observer
May 26.37851 17^{h}25^{m}09.^{s}40 +10^{o}15′28′′.7 18.1 LINEAR
26.42390 17 25 08.53 +10 15 57.2 18.7  _{”}
26.86019 17 25 00.95 +10 20 22.9 17.6 Tichá
26.87640 17 25 00.67 +10 20 32.4  _{”}
26.91351 17 24 59.91 +10 20 54.3  Kušnirák
26.92617 17 24 59.69 +10 21 01.9  _{”}
26.95036 17 24 59.19 +10 21 16.5  Galád
26.96775 17 24 58.88 +10 21 27.3 17.2  _{”}
2000 May 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green